The price of a pen mistakenly had its two digits reversed so that a customer was over charged 18 cents. If the sum of the digits is 16. Determine the correct price of the pen.


I've tried this problem multiple times and I just cunfuse my self even more.

Answer by ankor@dixie-net.com(22740)

The price of a pen mistakenly had its two digits reversed so that a customer was over charged 18 cents.
If the sum of the digits is 16. Determine the correct price of the pen.
:
let x = the 10 digit of the correct price
let y = the units
then
10x+y = the correct price
and
10y+x = the incorrect price which 18 cents more
:
"the sum of the digits is 16", therefore
x + y = 16
:
the equation for this situation:
10x+y = 10y+x - 18
Combine like terms
10x - x = 10y - y - 18
9x = 9y - 18
simplify, divide thru by 9, results:
x = y - 2
:
In the sum of the digits equation, replace x with (y-2)
y - 2 + y = 16
2y = 16 + 2
2y = 18
y = 9 is the units digit
then
16 - 9 = 7 is the 10's digit
:
The correct price: 79 cents
:
Check this
The incorrect price: 97 cents which is 18 cents more